Two small charged metal spheres at a distance of 0.50 m from each other attract each other with the force F = 5.0 μN. one has a charge of -3.5 nC. What charge does the other bullet have?


1
Expert's answer
2022-05-02T13:41:56-0400
"F=k\frac{q_1q_2}{r^2}"

"q_2=\frac{Fr^2}{kq_1}=\frac{5.0*10^{-6}*0.50^2}{9*10^9*3.5*10^{-9}}=4.0*10^{-8}\:\rm C"